# A variational discretization method for mean curvature flows by the Onsager principle

### v1

@@ 5.1 Derivation of the dynamic equation by the Onsager principle
~˙$\dot{\tilde{\mathcal{E}}}=-2\tilde{\Phi}_{h}\leq 0.$ (73)
@@
Similarly to the problems in the previous two section, we ignore the tangential velocity, the dynamic equation for a wetting problem can be described by
$\left\{\begin{array}[]{rcl}&\vec{v}=-\xi_{0}^{-1}\gamma(\kappa-\frac{\theta_{0}+\theta_{L}}{|\tilde{\Gamma}(t)|})\vec{n}\\ &v(0)=-\xi_{1}^{-1}\gamma(\cos\theta_{0}-\cos\theta_{Y})\\ &v(L)=\xi_{1}^{-1}\gamma(\cos\theta_{L}-\cos\theta_{Y}),\end{array}\right.$ (74)
where $\vec{v}$ denotes the velocity of the curve, $v(0)$ and $v(L)$ are respectively the horizontal velocity of the two endpoints, $\theta_{L}$ and $\theta_{0}$ denote the left and right contact angles of the curve with the bottom line, and $|\tilde{\Gamma}(t)|$ denotes the length of the curve.
@@ 5.2 Discretization of the dynamic equation by the Onsager principle
In this section, we aim to numercially solve the dynamic equation (65). Suppose that the curve $\tilde{\Gamma}(t)$ can be approximated by piecewise linear segments, as illustrated in Figure 4. Denote the nodes by
$\vec{x}_{1}(t),\vec{x}_{2}(t),...,\vec{x}_{n}(t).$ (75)
Where $\vec{x}_{i}(t)=(x_{i}^{(1)}(t),x_{i}^{(2)}(t))^{T}\ (1\leq i\leq n)$, the right endpoint $\vec{x}_{1}(t)=(x_{1}^{(1)}(t),0)^{T}$ and the left endpoint $\vec{x}_{n}(t)=(x_{n}^{(1)}(t),0)^{T}$. The line segments connecting adjacent nodes are denoted by $\tilde{\Gamma}_{i}(t)=\overline{\vec{x}_{i}(t)\vec{x}_{i+1}(t)}\ (1\leq i\leq n-1)$. Consequently, the approximate curve is given by $\tilde{\Gamma}_{h}(t)=\bigcup\limits_{i=1}^{n-1}\tilde{\Gamma}_{i}(t)$.
Figure 4: Selection of interpolation points
We will derive the dynamic equation for $\tilde{\Gamma}_{h}(t)$ by using the Onsager principle. Similar to that in Section 3.2, for a line segment $\tilde{\Gamma}_{i}(t)(1\leq i\leq n-1)$, the points on the segment can be expressed as $\vec{x}_{h}(t)=\mu[\vec{x}_{i+1}(t)-\vec{x}_{i}(t)]+\vec{x}_{i}(t)=(1-\mu)\vec{x}_{i}(t)+\mu\vec{x}_{i+1}(t),$ where $0\leq\mu=\frac{s}{|\vec{x}_{i+1}-\vec{x}_{i}|}\leq 1$ is a parameter. The velocity of each point $\vec{x}_{h}(t)$ is given by $\vec{v}_{h}(t)=(1-\mu)\dot{\vec{x}}_{i}(t)+\mu\dot{\vec{x}}_{i+1}(t).$ The unit tangent vector of $\tilde{\Gamma}_{i}(t)$ is calculated by $\vec{\tau}_{i}=\frac{\vec{x}_{i+1}(t)-\vec{x}_{i}(t)}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}.$ The outward unit normal vector is given by $\vec{n}_{i}=\frac{P[\vec{x}_{i+1}(t)-\vec{x}_{i}(t)]}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}.$
@@
$=\gamma\{\sum_{i=1}^{n-1}|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|-[x_{1}^{(1)}(t)-x_{n}^{(1)}(t)]\cos\theta_{Y}\}+C.$
The discrete dissipation funtion is calculated by
$\tilde{\Phi}_{h}$ $=\frac{1}{2}\int_{\tilde{\Gamma}_{h}(t)}v_{h,n}^{2}\text{d}s+\frac{1}{2}v_{h}(0)^{2}+\frac{1}{2}v_{h}(L)^{2}$ (77)
$=\frac{1}{2}\sum_{i=1}^{n-1}\int_{\tilde{\Gamma}_{i}(t)}[\vec{v}_{h}(t)\cdot\vec{n}_{i}]^{2}\text{d}s+\frac{1}{2}[\dot{x}_{1}^{(1)}(t)]^{2}+\frac{1}{2}[\dot{x}_{n}^{(1)}(t)]^{2}.$
@@
$v_{h,n}(0)=v_{h}(0)\sin\theta_{0}=\dot{x}_{1}^{(1)}(t)\sin\theta_{0},\quad v_{h,n}(L)=-v_{h}(L)\sin\theta_{L}=-\dot{x}_{n}^{(1)}(t)\sin\theta_{L},$ (78)
@@
$\sin\theta_{0}=\frac{x_{2}^{(2)}(t)}{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|},\quad\sin\theta_{L}=\frac{x_{n-1}^{(2)}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}.$ (79)
$\frac{\text{d}\tilde{S}_{h}}{\text{d}t}$ $=\int_{\tilde{\Gamma}_{h}(t)}v_{h,n}\text{d}s=\sum_{i=1}^{n-1}\int_{\tilde{\Gamma}_{i}(t)}\vec{v}_{h}(t)\cdot\vec{n}_{i}\text{d}s$ (80)
$\min_{\dot{\vec{x}}_{i}}\ \tilde{\mathcal{R}}_{h}:=\tilde{\Phi}_{h}+\dot{\tilde{\mathcal{E}}}_{h},$ (81)
@@
~$\tilde{\mathcal{L}}_{h}(\dot{\vec{x}}_{i},\lambda)=\tilde{\mathcal{R}}_{h}+\lambda\frac{\text{d}\tilde{S}_{h}}{\text{d}t}.$ (82)
@@
{~$\left\{\begin{array}[]{rcl}\frac{\partial\tilde{\mathcal{L}}_{h}}{\partial\dot{\vec{x}}_{i}}=0\\ \frac{\partial\tilde{\mathcal{L}}_{h}}{\partial\lambda}=0.\end{array}\right.$ (83)
@@
$\left\{\begin{array}[]{ll}\frac{\partial\tilde{\Phi}_{h}}{\partial\dot{x}_{1}^{(1)}}+\lambda\frac{\partial(\tilde{S}_{h})_{t}}{\partial\dot{x}_{1}^{(1)}}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{1}^{(1)}}\quad for\quad i=1\\ \frac{\partial\tilde{\Phi}_{h}}{\partial\dot{\vec{x}}_{i}}+\lambda\frac{\partial(\tilde{S}_{h})_{t}}{\partial\dot{\vec{x}}_{i}}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial\vec{x}_{i}}\qquad\!for\quad 2\leq i\leq n-1\\ \frac{\partial\tilde{\Phi}_{h}}{\partial\dot{x}_{n}^{(1)}}+\lambda\frac{\partial(\tilde{S}_{h})_{t}}{\partial\dot{x}_{n}^{(1)}}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{n}^{(1)}}\quad for\quad i=n\\ \frac{\text{d}\tilde{S}_{h}}{\text{d}t}=0.\end{array}\right.$ (84)
This is a system of differential-algebraic equations with respect to $\dot{x}_{1}^{(1)},\dot{\vec{x}}_{i}\ (2\leq i\leq n-1),\dot{x}_{n}^{(1)}$ and $\lambda$.
For convenience of representation, we add two equations $\dot{x}_{1}^{(2)}=0,\dot{x}_{n}^{(2)}=0$ to (84) and denote $\dot{X}=(\dot{\vec{x}}_{1}^{T},\dot{\vec{x}}_{2}^{T},...,\dot{\vec{x}}_{n}^{T},\lambda)^{T}$. Observe that $\tilde{\Phi}_{h}$ is a quadratic function with respect to $\dot{\vec{x}}_{i}\ (1\leq i\leq n)$. Then the system (84) can be rewritten as
@@
where the coefficient matrix $\tilde{A}$ is symmetric and the right-hand term $\tilde{G}=(\tilde{\vec{g}}_{1}^{T},\tilde{\vec{g}}_{2}^{T},...,\tilde{\vec{g}}_{n}^{T},\tilde{g}_{n+1})^{T}$. The explicit formulas for $\tilde{A}$ and $\tilde{G}$ are given in Appendix.
@@
Suppose $\dot{X}=(\dot{\vec{x}}_{1}^{T},\dot{\vec{x}}_{2}^{T},...,\dot{\vec{x}}_{n}^{T},\lambda)^{T}$ is the solution of the equation set (85), then we have
$\frac{\text{d}\tilde{\mathcal{E}}_{h}}{\text{d}t}=-2\tilde{\Phi}_{h}\leq 0,$ (86)
@@
~$\tilde{\mathcal{E}}_{h}^{\delta}=\tilde{\mathcal{E}}_{h}+\delta\sum_{i=1}^{n-2}\bigl{(}\frac{|\tilde{\Gamma}_{i}(t)|}{|\tilde{\Gamma}_{i+1}(t)|}-1\bigr{)}^{2}.$ (87)
@@ 6 Numerical experiments
$\left\{\begin{array}[]{rcl}\text{predictor}:{\tilde{X}}^{k+1}&=&X^{k}+\dot{X}^{k}\Delta t\\ \text{corrector}:X^{k+1}&=&X^{k}+\frac{(\dot{X}^{k}+\dot{{\tilde{X}}}^{k+1})}{2}\Delta t.\end{array}\right.$ (89)
where $X^{k}$ represents the position of the curve at time $t_{k}$. For the equation (26), $\dot{X}^{k}=A^{-1}(X^{k})G^{\delta}(X^{k})$ denotes the time derivative of $X^{k}$, and $\dot{\tilde{X}}^{k+1}=A^{-1}(\tilde{X}^{k+1})G^{\delta}(\tilde{X}^{k+1})$. For the equations (50) and (85), we use $\hat{A}$, $\hat{G}$ and $\tilde{A}$, $\tilde{G}$ respectively.
@@ 6.1.1 The evolution of a circular curve
$\Gamma(t):=\{\vec{x}(t,\alpha)=\rho(t)(\cos\alpha,\sin\alpha)^{T},\quad 0\leq\alpha\leq 2\pi\},$ (90)
where $\rho(t)$ denotes the radius of the circle at time $t$, and it satisfies the initial condition $\rho(0)=R_{0}$. By substituting the equation (90) into Equation (16), we have
@@
Solving the equation (91), we can obtain the evolution equation of the curve $\Gamma(t)$:
$\vec{x}(t,\alpha)=\sqrt{R_{0}^{2}-2t}(\cos\alpha,\sin\alpha)^{T},\quad 0\leq\alpha\leq 2\pi.$ (92)
@@
In our numerical simulations, the initial curve is taken as the unit circle. The circle is discretized uniformly with $\vec{x}_{i}=(\cos\frac{2\pi}{n}i,\sin\frac{2\pi}{n}i)^{T}$. The shape of the curve at some time steps are shown in Figure 5, while the change of the discrete energy function $\mathcal{E}_{h}(t)$ over time is shown in Figure 6. We can see that discrete curves are circular and the discrete energy decreases gradually.
@@
Since the explicit solution is known in this example (as shown in (92)), we can compute the errors of the numerical slutions. We denote the point on the discrete curve $\Gamma_{h}(t)$ by $\vec{x}_{h}(t,\alpha)$. We define the error as
$err:=\text{dist}(\vec{x}_{h}(t),\Gamma(t)).$ (93)
@@
$order:=\frac{\ln(\frac{err_{k}}{err_{k+1}})}{\ln(\frac{n_{k+1}}{n_{k}})}.$ (94)

### v3

@@ 5.1 Derivation of the dynamic equation by the Onsager principle
~˙$\dot{\tilde{\mathcal{E}}}=-2\tilde{\Phi}\leq 0.$
@@ 5.2 Discretization of the dynamic equation by the Onsager principle
In this section, we aim to numerically solve the dynamic equation (30d). Suppose that the curve $\tilde{\Gamma}(t)$ can be approximated by piecewise linear segments, as illustrated in Fig. 4. Denote the nodes by
$\vec{x}_{1}(t),\vec{x}_{2}(t),...,\vec{x}_{n}(t).$
Where $\vec{x}_{i}(t)=(x_{i}^{(1)}(t),x_{i}^{(2)}(t))^{\top}\ (1\leq i\leq n)$, the right endpoint $\vec{x}_{1}(t)=(x_{1}^{(1)}(t),0)^{\top}$ and the left endpoint $\vec{x}_{n}(t)=(x_{n}^{(1)}(t),0)^{\top}$. The line segments connecting adjacent nodes are denoted by $\tilde{\Gamma}_{i}(t)=\overline{\vec{x}_{i}(t)\vec{x}_{i+1}(t)}\ (1\leq i\leq n-1)$. Consequently, the approximate curve is given by $\tilde{\Gamma}_{h}(t)=\bigcup\limits_{i=1}^{n-1}\tilde{\Gamma}_{i}(t)$.
Figure 4: Selection of interpolation points.
We will derive the dynamic equation for $\tilde{\Gamma}_{h}(t)$ by using the Onsager principle. Similar to that in Section 3.2, for a line segment $\tilde{\Gamma}_{i}(t)\ (1\leq i\leq n-1)$, the points on the segment can be expressed as $\vec{x}_{h}(t)=(1-\mu)\vec{x}_{i}(t)+\mu\vec{x}_{i+1}(t),$ where $0\leq\mu=\frac{s}{|\vec{x}_{i+1}-\vec{x}_{i}|}\leq 1$ is a parameter and $s$ is the arc length parameter of $\tilde{\Gamma}_{i}(t)$. The velocity of each point $\vec{x}_{h}(t)$ is given by $\vec{v}_{h}(t)=(1-\mu)\dot{\vec{x}}_{i}(t)+\mu\dot{\vec{x}}_{i+1}(t).$ The unit tangent vector of $\tilde{\Gamma}_{i}(t)$ is calculated by $\vec{\tau}_{i}=\frac{\vec{x}_{i+1}(t)-\vec{x}_{i}(t)}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}.$ The outward unit normal vector is given by $\vec{n}_{i}=\frac{P[\vec{x}_{i+1}(t)-\vec{x}_{i}(t)]}{|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|}.$
@@
$=\gamma\left\{\sum_{i=1}^{n-1}|\vec{x}_{i+1}(t)-\vec{x}_{i}(t)|-[x_{1}^{(1)}(t)-x_{n}^{(1)}(t)]\cos\theta_{Y}\right\}+C.$
The discrete dissipation function is calculated by
$\tilde{\Phi}_{h}$ $=\frac{\xi_{0}}{2}\int_{\tilde{\Gamma}_{h}(t)}|\vec{v}_{h}|^{2}\text{d}s+\frac{\xi_{1}}{2}v_{h,0}^{2}+\frac{\xi_{1}}{2}v_{h,L}^{2}$
$=\frac{\xi_{0}}{2}\sum_{i=1}^{n-1}\int_{\tilde{\Gamma}_{i}(t)}[(1-\mu)\dot{\vec{x}}_{i}(t)+\mu\dot{\vec{x}}_{i+1}(t)]^{2}\text{d}s+\frac{\xi_{1}}{2}[\dot{x}_{1}^{(1)}(t)]^{2}+\frac{\xi_{1}}{2}[\dot{x}_{n}^{(1)}(t)]^{2}.$
$\frac{\text{d}\tilde{S}_{h}}{\text{d}t}$ $=\int_{\tilde{\Gamma}_{h}(t)}v_{h,n}\text{d}s=\sum_{i=1}^{n-1}\int_{\tilde{\Gamma}_{i}(t)}\vec{v}_{h}(t)\cdot\vec{n}_{i}\text{d}s.$
@@
$\dot{\vec{x}}_{1}(t)\cdot\vec{n}_{1}=\dot{x}_{1}^{(1)}(t)\sin\theta_{0},\quad\dot{\vec{x}}_{n}(t)\cdot\vec{n}_{n-1}=-\dot{x}_{n}^{(1)}(t)\sin\theta_{L},$
@@
$\sin\theta_{0}=\frac{x_{2}^{(2)}(t)}{|\vec{x}_{2}(t)-\vec{x}_{1}(t)|},\quad\sin\theta_{L}=\frac{x_{n-1}^{(2)}(t)}{|\vec{x}_{n}(t)-\vec{x}_{n-1}(t)|}.$
By using the Onsager principle, we minimize the discrete Rayleighian with respect to $\dot{\vec{x}}_{i}(2\leq i\leq n-1),\dot{x}_{1}^{(1)},\dot{x}_{n}^{(1)}$ under the area conservation constraint, i.e.
$\min_{\dot{\vec{x}}_{i},\dot{x}_{1}^{(1)},\dot{x}_{n}^{(1)}}\ \tilde{\mathcal{R}}_{h}:=\tilde{\Phi}_{h}+\dot{\tilde{\mathcal{E}}}_{h},$ (36)
@@
~$\tilde{\mathcal{L}}_{h}(\dot{\vec{x}}_{i},\dot{x}_{1}^{(1)},\dot{x}_{n}^{(1)},\lambda)=\tilde{\mathcal{R}}_{h}+\lambda\frac{\text{d}\tilde{S}_{h}}{\text{d}t}.$
@@
{~$\left\{\begin{array}[]{rcl}\frac{\partial\tilde{\mathcal{L}}_{h}}{\partial\dot{\vec{x}}_{i}}=0,\vspace{1ex}\\ \frac{\partial\tilde{\mathcal{L}}_{h}}{\partial\dot{x}_{1}^{(1)}}=0,\vspace{1ex}\\ \frac{\partial\tilde{\mathcal{L}}_{h}}{\partial\dot{x}_{n}^{(1)}}=0,\vspace{1ex}\\ \frac{\partial\tilde{\mathcal{L}}_{h}}{\partial\lambda}=0.\end{array}\right.$
@@
$\left\{\begin{array}[]{ll}\frac{\partial\tilde{\Phi}_{h}}{\partial\dot{\vec{x}}_{i}}+\lambda\frac{\partial(\tilde{S}_{h})_{t}}{\partial\dot{\vec{x}}_{i}}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial\vec{x}_{i}},\quad 2\leq i\leq n-1,\vspace{1ex}\\ \frac{\partial\tilde{\Phi}_{h}}{\partial\dot{x}_{1}^{(1)}}+\lambda\frac{\partial(\tilde{S}_{h})_{t}}{\partial\dot{x}_{1}^{(1)}}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{1}^{(1)}},\vspace{1ex}\\ \frac{\partial\tilde{\Phi}_{h}}{\partial\dot{x}_{n}^{(1)}}+\lambda\frac{\partial(\tilde{S}_{h})_{t}}{\partial\dot{x}_{n}^{(1)}}=-\frac{\partial\tilde{\mathcal{E}}_{h}}{\partial x_{n}^{(1)}},\vspace{1ex}\\ \frac{\text{d}\tilde{S}_{h}}{\text{d}t}=0.\end{array}\right.$ (37)
This is a system of differential-algebraic equations with respect to $\dot{\vec{x}}_{i}\ (2\leq i\leq n-1),\dot{x}_{1}^{(1)},\dot{x}_{n}^{(1)}$ and $\lambda$.
For convenience of representation, we add two equations $\dot{x}_{1}^{(2)}=0,\dot{x}_{n}^{(2)}=0$ to (37) and denote $\dot{X}=(\dot{\vec{x}}_{1}^{\top},\dot{\vec{x}}_{2}^{\top},...,\dot{\vec{x}}_{n}^{\top},\lambda)^{\top}$. Observe that $\tilde{\Phi}_{h}$ is a quadratic function with respect to $\dot{\vec{x}}_{i}\ (1\leq i\leq n)$. Then the system (37) can be rewritten as
@@
where $\tilde{A}$ is a symmetric matrix and $\tilde{G}=(\tilde{\vec{g}}_{1}^{\top},\tilde{\vec{g}}_{2}^{\top},...,\tilde{\vec{g}}_{n}^{\top},\tilde{g}_{n+1})^{\top}$. The explicit formulas for $\tilde{A}$ and $\tilde{G}$ are given in Appendix.
@@
Suppose $\dot{X}=(\dot{\vec{x}}_{1}^{\top},\dot{\vec{x}}_{2}^{\top},...,\dot{\vec{x}}_{n}^{\top},\lambda)^{\top}$ is the solution of Eq. (38), then we have
$\frac{\text{d}\tilde{\mathcal{E}}_{h}}{\text{d}t}=-2\tilde{\Phi}_{h}\leq 0,$
@@
~$\tilde{\mathcal{E}}_{h}^{\delta}=\tilde{\mathcal{E}}_{h}+\delta\sum_{i=1}^{n-2}\left(\frac{|\tilde{\Gamma}_{i}(t)|}{|\tilde{\Gamma}_{i+1}(t)|}-1\right)^{2}.$
@@ 6 Numerical experiments
$\left\{\begin{array}[]{rcl}\text{predictor}:{\bar{\bar{X}}}^{k+1}&=&X^{k}+\dot{X}^{k}\Delta t,\vspace{1ex}\\ \text{corrector}:X^{k+1}&=&X^{k}+\frac{(\dot{X}^{k}+\dot{{\bar{\bar{X}}}}^{k+1})}{2}\Delta t.\end{array}\right.$ (40)
Where $X^{k}$ represents the position of the curve at time $t_{k}$. For Eq. (15), $\dot{X}^{k}=A^{-1}(X^{k})G^{\delta}(X^{k})$ denotes the time derivative of $X^{k}$, and $\dot{\bar{\bar{X}}}^{k+1}=A^{-1}(\bar{\bar{X}}^{k+1})G^{\delta}(\bar{\bar{X}}^{k+1})$. For Eqs. (25) and (39), we use $\hat{A}$, $\hat{G}^{\delta}$ and $\tilde{A}$, $\tilde{G}^{\delta}$ respectively.
@@ 6.1.1 The evolution of a circular curve
$\Gamma(t):=\{\vec{x}(\alpha,t)=\rho(t)(\cos\alpha,\sin\alpha)^{\top},\quad 0\leq\alpha\leq 2\pi\},$ (41)
where $\rho(t)$ denotes the radius of the circle at time $t$, and it satisfies the initial condition $\rho(0)=R_{0}$. By substituting Eq. (41) into Eq. (9), we have
@@
Solving Eq. (42), we can obtain the evolution equation of the curve $\Gamma(t)$
$\vec{x}(\alpha,t)=\sqrt{R_{0}^{2}-2t}(\cos\alpha,\sin\alpha)^{\top},\quad 0\leq\alpha\leq 2\pi.$ (43)
@@
In our numerical simulations, the initial curve is taken as the unit circle. The circle is discretized uniformly with $\vec{x}_{i}=(\cos\frac{2\pi}{n}i,\sin\frac{2\pi}{n}i)^{\top}$. The shape of the curve at some time steps are shown in Fig. 5, while the change of the discrete energy function $\mathcal{E}_{h}(t)$ over time is shown in Fig. 6. We can see that discrete curves are circular and the discrete energy decreases gradually.
@@
Since the explicit solution is known in this example (as shown in (43)), we can compute the errors of the numerical solutions. We denote the point on the discrete curve $\Gamma_{h}(t)$ by $\vec{x}_{h}(\alpha,t)$. We define the error as
$err_{1}:=\max_{\vec{x}_{h}\in\Gamma_{h}(t)}\text{dist}(\vec{x}_{h},\Gamma(t)).$ (44)
@@
$order:=\frac{\ln(err_{1}^{(k)}/err_{1}^{(k+1)})}{\ln(n_{k+1}/n_{k})}.$ (45)